# Eigenstate thermalization to non-monotonic distributions in strongly-interacting chaotic lattice gases

### v2

@@ Appendix B One-dimensional Bose-Hubbard model
where $l=1,...,L$ is the site index, $J_{lm}=J_{ml}^{*}$ is the hopping matrix coupling sites $l$ and $m$, $V$ is the on-site interaction strength, $\hat{n}_{l}=\hat{b}_{l}^{\dagger}\hat{b}_{l}$ is the number of bosons at site $l$, and $b_{l}$ are bosonic particle annihilation operators. Throughout the manuscript we have used the Bose-Hubbard (BH) configuration $J_{l\neq m}=\delta_{l,m\pm 1}$ with hard wall boundaries, i.e. a linear chain of $L$ sites. For this configuration, the dynamical behavior of the system, e.g. its degree of chaoticity, is set by the dimensionless interaction parameter $u=VN$. In order to remove the remaining parity symmetry and increase chaoticity, we have introduced a weak random ’disorder’ on-site potential $J_{l,l}={\rm rnd}[-0.05,0.05]$.
@@ Appendix C Chaotic properties
The degree of chaoticity of a quantum system can be deduced from its level spacing statistics. One measure of the transition from the Poissonian statistics of integrable systems to the Wigner-Dyson statistics of completely chaotic systems is the ratio of consecutive level spacings [29]
@@
averaged over the entire spectrum or over a pertinent energy shell. The value $\left\langle r\right\rangle=2\ln 2-1\approx 0.38629$ is indicative of Poissonian statistics, whereas $\left\langle r\right\rangle=4-2\sqrt{3}\approx 0.53590$ is obtained for Wigner-Dyson GOE statistics [30]. In Fig. 4 we present this measure as a function of the interaction strength for our model systems. The 1D BH system is integrable at weak interaction due to its near-separability and at strong interaction due to macroscopic self-trapping where site occupations become integrals of motion. By contrast, the 2D FH system does not return to integrability at high interaction strength.
@@
Figure 5: Ratio of eigenstate-to-eigenstate fluctuation variances for the non-integrable to ones for the integrable systems eigenstates vs. interaction strength for: (a) the FH model averaged over the microcanonical shell with the mean energy 00; (b) the 1D BH model with the mean shell energy $-5.14$.
The variances are presented in Fig. 5. Due to the large number of the orbitals in the FH model, we consider cumulative observables: the total occupations of orbitals with $m_{x}<L_{x}/2$ and any $m_{y}$ in Eq. (A.3) [low x in Fig. 5(a)], with $m_{y}<L_{y}/2$ and any $m_{x}$ (low y), with $m_{x}<L_{x}/2$ and $m_{y}<L_{y}/2$ (low xy), with even $m_{x}$ and any $m_{y}$ (even x), with even $m_{y}$ and any $m_{x}$ (even y), and with even $m_{x}$ and $m_{y}$ (even xy). We also consider the hopping energies in the $x$
$-\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\sum_{\delta_{x}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}+\delta_{x}l_{y}}$
@@
$-\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\sum_{\delta_{y}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}+\delta_{y}}$
@@
Figure 6: Local density of states, averaged over the orbitals for: (a) the FH model with $V=10$ at $E=-5.2$; (b) the FH model with $V=10$ at $E=38.4$; (c) the 1D BH model with $V=3$ at $E=-5.14$.
@@ Appendix D Local density of states
$\bar{W}(E,\bar{E}_{n})=\frac{\Delta_{n}}{\Delta_{\mathrm{MC}}}\sum_{n\in\mathrm{MC}(\bar{E}_{n})}W(E,E_{n})$
and the averaged LDOS divided by $\Delta_{n}$ in a comparison with the Gaussian profiles.
In the centre of the spectrum, both $\bar{W}(E,\bar{E}_{n})$ and $\bar{W}(E,\bar{E}_{n})/\Delta_{n}$ can be approximated by Gaussian profiles (see Fig. 6(a) and (c)). However, near the spectrum boundaries only $\bar{W}(E,\bar{E}_{n})/\Delta_{n}$ has a Gaussian shape, but $\bar{W}(E,\bar{E}_{n})$ does not (see Fig. 6(b)).

### v4

@@ Appendix B One-dimensional Bose-Hubbard model
where l=1,,Ll=1,...,L is the site index, $J_{lm}=J_{ml}^{*}$ is the hopping matrix coupling sites $l$ and $m$, $V$ is the on-site interaction strength, $\hat{n}_{l}=\hat{b}_{l}^{\dagger}\hat{b}_{l}$ is the number of bosons at site $l$, and $\hat{b}_{l}$ are bosonic particle annihilation operators. Throughout the manuscript we have used the Bose-Hubbard (BH) configuration $J_{l\neq m}=\delta_{l,m\pm 1}$ with hard wall boundaries, i.e. a linear chain of $L$ sites. For this configuration, the dynamical behavior of the system, e.g. its degree of chaoticity, is set by the dimensionless interaction parameter $u=VN$. In order to remove the remaining parity symmetry and increase chaoticity, we have introduced a weak random ’disorder’ on-site potential $J_{l,l}={\rm rnd}[-0.05,0.05]$.
@@ Appendix C Chaotic properties
The degree of chaoticity of a quantum system can be deduced from its level spacing statistics. One measure of the transition from the Poissonian statistics of integrable systems to the Wigner-Dyson statistics of completely chaotic systems is the ratio of consecutive level spacings
@@
averaged over the entire spectrum or over a pertinent energy shell. This criterion has been introduced in [30], and is very widely used as a clear evidence of integrability-chaos transition. The value $\left\langle r\right\rangle=2\ln 2-1\approx 0.38629$ is indicative of Poissonian statistics, whereas $\left\langle r\right\rangle=4-2\sqrt{3}\approx 0.53590$ is obtained for Wigner-Dyson GOE statistics [31]. In Fig. 4 we present this measure as a function of the interaction strength for our model systems. The 1D BH system is integrable at weak interaction due to its near-separability and at strong interaction due to macroscopic self-trapping where site occupations become integrals of motion. In contrast, the 2D FH system does not return to integrability at high interaction strength.
@@
Figure 5: Ratio of eigenstate-to-eigenstate fluctuation variances for the non-integrable to ones for the integrable systems eigenstates vs. interaction strength for: (a) the FH model averaged over the microcanonical shell with the mean energy 0; (b) the 1D BH model with the mean shell energy 5.14-5.14.
Figure 6: Local density of states, averaged over the orbitals for: (a) the FH model with $V=10$ at E=5.2E=-5.2; (b) the FH model with $V=10$ at $E=38.4$; (c) the 1D BH model with $V=3$ at E=5.14E=-5.14.
The variances are presented in Fig. 5. Due to the large number of the orbitals in the FH model, we consider cumulative observables: the total occupations of orbitals with $m_{x}<L_{x}/2$ and any $m_{y}$ in Eq. (LABEL:eq:pxy) [low x in Fig. 5(a)], with $m_{y}<L_{y}/2$ and any $m_{x}$ (low y), with $m_{x}<L_{x}/2$ and $m_{y}<L_{y}/2$ (low xy), with even $m_{x}$ and any $m_{y}$ (even x), with even $m_{y}$ and any $m_{x}$ (even y), and with even $m_{x}$ and $m_{y}$ (even xy). We also consider the hopping energies in the $x$
$\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\sum_{\delta_{x}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}+\delta_{x}l_{y}}$
@@
$\sum_{l_{x}=1}^{L_{x}}\sum_{l_{y}=1}^{L_{y}}\sum_{\delta_{y}=\pm 1}\hat{a}_{l_{x}l_{y}}^{\dagger}\hat{a}_{l_{x}l_{y}+\delta_{y}}$
@@ Appendix D Local density of states
$\bar{W}(E,\bar{E}_{n})=\frac{\Delta_{n}(E_{n})}{\Delta_{\mathrm{MC}}}\sum_{n\in\mathrm{MC}(\bar{E}_{n})}W(E,E_{n})$
and the averaged LDOS divided by $\Delta_{n}(E_{n})$ in a comparison with the Gaussian profiles.
In the centre of the spectrum, both $\bar{W}(E,\bar{E}_{n})$ and $\bar{W}(E,\bar{E}_{n})/\Delta_{n}(E_{n})$ can be approximated by Gaussian profiles (see Figs. 6(a) and (c)). However, near the spectrum boundaries, where the spectrum is substantially inhomogeneous, only $\bar{W}(E,\bar{E}_{n})/\Delta_{n}(E_{n})$ has a Gaussian shape, but $\bar{W}(E,\bar{E}_{n})$ does not (see Fig. 6(b)).